A block of mass 5 kg is on a tabletop (which we will assume for the moment is frictionless) A string tied to the block runs over a pulley and is tied to a 2 kg block hanging below the pulley
a) How much gravity force acts on the 5 kg block? Does this cause any motion?
b) How much force acts on the 2 kg block? Does this cause any motion?
c) How much force actually causes things to move?
d) How much mass is being moved ?
e) How much acceleration occurs?
f) How much force must be acting in the string where it is tied to the 5 kg block?
g) Now repeat the whole thing assuming that a friction force of 5 N. Acts to slow everything down.
1
Expert's answer
2019-11-18T10:44:47-0500

a)Gravity force acts on 5kg block "G_1=m_1g=5\cdot 10=50 (N)". No moving in this case.

b)Gravity force acts on 5kg block "G_2=m_2g=2\cdot 10 = 20 (N)". No moving in this case.

c)Force actually causes things to move "F=G_1-G_2=30(N)"

d)Mass "G_3=30; m_3=\frac{G_3}{g}=\frac{30}{10}=3 (kg)"

e) Acceleration - 0

f)Force must be acting in the string "T=m_2g=20(N)"

d)-----

a)Gravity force acts on 5kg block "G_1=m_1g=5\cdot 10=50 (N)" No moving in this case.

b)Gravity force acts on 5kg block "G_2=m_2g=2\cdot 10 = 20 (N)" . No moving in this case.

c)Force actually causes things to move"F=G_1-G_2+F_f=35(N)"

d)Mass "G_3=35(N); m_3=\frac{G_3}{g}=\frac{35}{10}=3.5 (kg)"

e) Acceleration - 0

f)Force must be acting in the string "T=m_2g-F_f=15(N)"